OpenStudy (anonymous):

Find the surface area of a sphere with a great circle that has an area of 49 π square miles. A. 225π mi2 B. 196 π mi2 C. 126 π mi2 D. 343 π mi2

OpenStudy (anonymous):

Okay so, Area of a circle = pir^2 So pir^2 = 49pi Pi pi get cancelled, so r^2 = 49 Or r = 7 mi So the surface area of a sphere is 4pir^2 Therefore, 4pir^2 = 4pi(49) 4*49=196pi mi^2. The answer is B. :)

OpenStudy (anonymous):

THANKS!